# Applying statistical learning theory to deep learning

### v1

@@ 2.4 What we don’t yet fully understand: benign overfitting
Connection with double descent.
@@
The fact that we get an increase followed by a decrease is not surprising. What is surprising is that we get good generalization even though we insist on 0 training error in a fairly noisy situation. The main experiment that Misha and coauthors did that probably convinced many people is from [5], in which high levels of noise were added to synthetic data. When perfectly fitting this noisy data using different methods (all methods get 0 training error), the test error is substantially below the null risk. In particular, the phenomenon we are seeing is that we are perfectly fitting the noise, but this overfitting is not harmful, as in Figure 5. Rather, the overfitting is benign. We are fitting the noise in a way that has a kind of measure-0 effect; fitting the noise does not ruin the fit in other places. We are maybe not gaining anything from fitting this noise, but it doesn’t hurt us either (it’s benign).
Open Questions
In what situations is overfitting harmful, and it what situations is it benign? In least squares, we can get both kinds of behaviors in a now-predictable way (the result of research in the past 2-3 years). We know that it relates to a measure of effective rank. Characterizing more generally (beyond least squares) when overfitting is harmful vs. benign is a big challenge that we have now, which is fairly different from how we used to think about overfitting.
In some sense, the most practical implication of benign overfitting is that we don’t have to worry too much about selecting the right value of the regularization parameter $\lambda$; we have a whole regime of good values of $\lambda$. In many cases in practice, though, we see something in between benign and harmful overfitting (but much closer to benign). This matches what we see empirically: that adding a bit of regularization can be helpful by a bit, but we can still have good performance without explicit regularization.
Figure 4: Trading off training loss and complexity.
@@
2.5 Summary
What fits our understanding and what doesn’t?
In linear models, which is the only case we really understand, we can see how we can generalize even when we are able to fit random labels. We can understand how the complexity control comes purely from the optimization algorithm and how we can generalize better with bigger and bigger models, since they provide a better and better approximation to what we really want. Therefore, to fit each case within our classic understanding, we “just” need to ask: 1. What is the complexity measure? 2. How is it minimized by the optimization algorithm? 3. How does minimizing it ensure good generalization? Thus, the only thing that truly goes beyond our standard understanding is the phenomenon of benign overfitting.
@@ 3.1 Learning and optimization for convex problems
Recall that the goal of supervised learning is, for given input and output spaces $\mathcal{X},\mathcal{Y}$, to find a predictor function $h_{w}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$, parametrized by a vector $w$, with low population error defined by:
@@
Here, the term $\Psi(w)$ is the regularizer or penalty, while the parameter $\lambda>0$ tunes the regularization strength. The purpose of this term is mainly to prevent overfitting. Intuitively, a well chosen $\Psi$ will increase if a corresponding complexity measure of the model increases. Typical examples include the $\ell_{2}$ norm, enforcing regularity at the functional level, or the $\ell_{1}$ norm, inducing sparsity at the level of the parameters $w$. Equivalently, (11) can be written as
@@
where $B$ depends on the regularization coefficient $\lambda$. The main two sources of error that need to be controlled in empirical risk minimization are the optimization and generalization error. The latter is handled using uniform convergence tools to control the convergence rate of the empirical risk towards the population one, when the parameters are constrained to the sublevel sets defined by the penalty.
@@
ensuring that the predictor $\hat{w}$ is a good proxy for the true minimizer. The standard way to achieve uniform control of the deviation between $\hat{L}(w)$ and $L(w)$ is to quantify the complexity of the hypothesis class $\mathcal{H}$ and regularity of the loss function, and to relate these quantities to the required sample complexity. A useful complexity measure for hypothesis classes is the Rademacher complexity, defined in its empirical form as
@@
$L(h)-\hat{L}(h)\leq 2\mathcal{R}(l\circ\mathcal{H}\circ S)+c\sqrt{\frac{2/\delta}{m}}$ (16)
@@
$\mathbb{E}[\nabla l(\langle\bar{w}^{k},x_{i}\rangle,y_{i})]=\nabla\hat{L}(\bar{w}^{k}).$
@@
Thus the error magnitude of the SGD and approximation error is the same. This means that we need to do at most $m$ iteration of SGD because when $m>T$, the dominant term in the previous bound becomes $O(\sqrt{\frac{B^{2}}{m}})$.
@@ 3.2 Stochastic optimization
Stochastic optimization problem is written as
$\min_{w\in{\mathcal{W}}}F(w)=\mathbb{E}_{z\sim{\mathcal{D}}}[f(w,z)]$ (20)
@@
$\min_{h}F(h)=\mathbb{E}_{z\sim{\mathcal{D}}}[f(h,z)],$
using the data samples $z_{i}\sim{\mathcal{D}}$, for $i=1,\ldots,m$, where $h$ is a mathematical object adapted to the particular model. Here is a short list of examples.
@@
    In both cases $L(w^{*})$ is the value of the objective at the optimal point in the class: $L(w^{*})=\min_{\|w\|_{2}\leq B}L(w)$.
@@
Although we mentioned the effect of the step-size on the regularization in the direct approach, it is still not clear why we observe this phenomenon. Let us look back at the standard GD. The gradient descent minimizes the norm $\|w\|_{2}$. Indeed. At each step of the gradient descent we minimize its linear approximation given by the gradient:
@@
The latter is linear function and thus its minimum is at infinity. Also, on the other hand, the linear approximation is not valid when we get far from the iterate $w^{(t)}$. That is why we add a square regularizer $\|w-w^{t}\|_{2}^{2}/2\eta$. We obtain
@@
In order to better understand this for the SGD, let us first introduce the notion of stability.
@@ 3.3 Stability
Define $\widehat{F}(w):=\frac{1}{m}\sum_{i=1}^{m}f_{i}(w)$. If $\tilde{w}$ is symmetric and $\beta(m)$ is stable then
@@ 3.4 Strong convexity
The function $\Psi:{\mathcal{W}}\rightarrow\mathbb{R}$ is $\alpha$-strongly convex w.r.t. to a norm $\|w\|$ if
@@
Strong convexity essentially means that the function is bounded below by a quadratic function. This property depends on the norm unlike the convexity which can be defined on a vector space. Let us apply the (23) for the optimum point $w_{0}:=\operatorname*{arg\,min}_{w}\Psi(w)$. Since $\nabla\Psi(w_{0})=0$ we get the following for every $w$:
@@
The latter means that the difference between the function value at any point and the optimal one is proportional to the distance from the optimal point. Let us now establish the connection of strong convexity and stability. For the dataset $S=\{z_{1},\ldots,z_{m}\}$ we define the regularized ERM (RERM) as follows:
@@
Here $\Psi$ is an $\alpha$-strongly convex function and $\widehat{F}$ is the empirical mean corresponding to the dataset $S$. We recall the definition of Lipschitz continuity.
@@
One may notice that Lipschitz continuity implies some kind of “stability”, as it essentially means that small perturbation of the argument will not change the function value drastically. As for strong convexity, the Lipschitzness also depends on the norm. This yields that $\|\nabla_{w}f(w,z)\|_{*}\leq G$, where $*\|\cdot\|_{*}$ is the dual norm. We will assume that the norm is the same for both properties.
@@
In the rest of the lecture, without loss of generality, we may assume that the regularization function $\Psi$ is $1$-strongly convex. This is easily achieved by tuning the parameter $\lambda$ accordingly. Theorem 3.1 yields the following
@@
where the last line is obtained by choosing $\lambda=\sqrt{\frac{2G^{2}}{\alpha m\sup_{w}\Psi(w)}}$. In particular, using the constraint $\Psi(w)\leq B$, the last inequality can be rewritten as
@@
The last inequality is obtain by optimizing the right-hand side w.r.t. to the parameter $\lambda$. Let us now look back at the Risk minimization problem in the convex case:
@@
${\mathcal{W}}$ is assumed to be convex. If ${\sf loss}(\hat{y},y)$ is convex in $\hat{y}$, then the problem is convex. For a non-trivial loss e.g. $\textsf{loss}(h_{w}(x),y)$ is convex in $w$ only when $h_{w}(x)=\langle w,\phi(x)\rangle$. In this setting the Lipschitz continuity goes as follows. Assume that the loss function $\sf loss(y,y^{\prime})$ is $g$-Lipschitz continuous w.r.t. $y$. Then
@@
The stochastic optimization resembles the online learning problem. The optimizer provides $w_{i}$. We give it to the adversary to compute $f(w_{i},z_{i})$ and then use the value of $f(w_{i},z_{i})$ to compute $w_{i+1}$.
@@
$\hat{w}_{m}(z_{1},\ldots,z_{m-1})=\arg\min_{w\in{\mathcal{W}}}\sum_{i=1}^{m}f(w,z_{i})+\lambda_{t}\Psi(w).$
This algorithm however, does not resemble to the one-pass SGD algorithm (see also equation (22)):
@@
$\hat{w}_{m}^{\lambda}(z_{1},\ldots,z_{m-1})=\arg\min_{w\in{\mathcal{W}}}\frac{1}{m}\left\langle\sum_{i=1}^{m}\nabla f(w_{i},z_{i}),w\right\rangle+\lambda_{t}\Psi(w).$ (38)
@@ 3.5 Mirror descent
Figure 8: Graphical representation of the Bregman divergence.
@@
where $\Pi_{\Psi}^{{\mathcal{W}}}(w)=\min_{w^{\prime}\in{\mathcal{W}}}D_{\Psi}(w^{\prime}\mid w)$ is a projection set on $\mathcal{W}$ with respect to the Bregman distance. One may easily verify that if $\Psi$ is a quadratic function, then its Bregman divergence is also quadratic, and we recover the standard gradient descent algorithm, assuming the constraint set $\mathcal{W}$ is the domain of definition of $\Psi$. Indeed, taking $\Psi(x)=\frac{1}{2}\lVert x\rVert_{2}^{2}$, we obtain
@@
Figure 9: Suppose that the hypothesis class ${\mathcal{W}}$ is a subset of ${\mathcal{B}}$, where ${\mathcal{B}}$ is a Banach space. As we have seen above, the gradients live in the dual space ${\mathcal{B}}^{*}$. Then each iteration of the mirror descent consists of four steps. First we take the map $\nabla\Psi:{\mathcal{B}}\rightarrow{\mathcal{B}}^{*}$. Then we do a step in ${\mathcal{B}}^{*}$ in gradient direction $\nabla f$. At the third step, we take $\Psi^{-1}$, which maps ${\mathcal{B}}^{*}$ to ${\mathcal{B}}$. The final step is the projection of this point to the hypothesis set ${\mathcal{W}}$.
@@
3.6 Summary
We presented guarantees for general and stochastic optimization under simple assumptions on the function being optimized and combined these with generalization guarantees to understand how these interact (in particular, is optimization or generalization the “harder” thing to do, in terms of sample complexity?). We studied stability and used it to motivate mirror descent, an abstraction that covers both one-pass SGD and linearized FTRL. Finally, we saw two different ways to see mirror descent. The first is based on the link between the primal and the dual spaces. However, it is less transparent as the link function $\Psi$ does not include information about the reality or model complexity. The other way, is the connection with the linearized FTRL. Here, $\Psi$ acts as a regularizer which naturally describes the link function as a complexity measure.
@@ 4 Lecture 4 : Mirror descent and implicit bias of descent algorithms
To recap, we talked last time about learning being a stochastic optimization problem. We are going to use z to get unbiased estimators of the gradient of the population loss. For learning, we consider an objective of the form:
@@
so that we have convexity with respect to $w\,.$ This is suitably generic because any data set is realizable in this form: Let $\phi$ map $x$ to an indicator about its identity, and $w$ selects the appropriate label.
There are two reasons to study convex optimization in a deep learning school: one is because it relates to mirror descent, and the second is that it allows us to talk about the geometry of the optimization problem and therefore the inductive bias.

### v2

@@ 2.4 A more recent form of implicit bias : benign overfitting
Connection with the so-called double descent.
@@
The fact that we get an increase followed by a decrease is not surprising. What is surprising is that we get good generalization even though we insist on 0 training error (a solution that interpolates the training set) in a fairly noisy situation. The main experiment showcasing this behaviour from [9], in which high levels of noise were added to synthetic data. When perfectly fitting this noisy data using different methods (all methods get 0 training error), the test error is substantially below the null risk. In particular, the phenomenon we are seeing is that we are perfectly fitting the noise, but this overfitting is not harmful, as in Figure 5. Rather, the overfitting is benign. We are fitting the noise in a way that has a kind of measure-0 effect; fitting the noise does not ruin the fit in other places. We are maybe not gaining anything from fitting this noise, but it doesn’t hurt us either (it’s benign).
Open questions regarding benign overfitting
In what situations is overfitting harmful, and in what situations is it benign? In least squares, we can get both kinds of behaviors in a now-predictable way (the result of research in the past 2-3 years, see e.g. [24, 6]). We know that it relates to a measure of effective rank at the level of the covariance matrix of the design matrix. Characterizing more generally (beyond least squares) when overfitting is harmful or benign is a big challenge that is yet to be solved, and is fairly different from how we used to think about overfitting.
In some sense, the most practical implication of benign overfitting is that we don’t have to worry too much about selecting the right value of the regularization parameter $\lambda$; we have a whole regime of good values of $\lambda$, even in noisy cases. In many cases in practice, though, we see something in between benign and harmful overfitting, although much closer to benign. This matches what we see empirically: that adding a bit of regularization can be helpful by a bit, but we can still have good performance without explicit regularization.
@@
2.5 A more general look at the implicit bias in optimization
From now on and in the remaining lectures, we will focus on understanding the implicit bias coming from optimization and leave aside the benign overfitting phenomenon. The examples above allow us to establish the following guidelines to understand implicit bias in optimization :
  • what complexity measure is actually being minimized by a given algorithm? This will depend on the choice of the algorithm but also, as we have seen, on the chosen architecture for our model.
  • how do the parameters of the algorithm (initialization scale, step size, etc …) affect the minimization of this complexity measure?
  • How do low values of this complexity measure ensure good generalization?
The goal of the remaining lectures will be to provide a framework and method to answer these questions in simple cases. In particular, we will rely on the generic formulation of the mirror descent algorithm, which will allow us to understand what cost function is implicitly being minimized for a given choice of architecture and algorithm parameters, as well as capture the geometry of the function space in which our algorithm is implicitly searching. Before doing so, let us provide some reminders on optimization in the context of supervised learning, which will be the subject of the next lecture.
@@ 3.1 Learning and optimization for convex problems
Recall that the objective of supervised learning is, for given input and output spaces $\mathcal{X},\mathcal{Y}$, to find a predictor function $h_{w}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$, parametrized by a vector $w$, with low population error defined by:
@@
Here, the term $\Psi(w)$ is the regularizer or penalty, while the parameter $\lambda>0$ tunes the regularization strength. The purpose of this term is mainly to prevent overfitting. Intuitively, a well chosen $\Psi$ will increase if a corresponding complexity measure of the model increases. Typical examples include the $\ell_{2}$ norm, enforcing regularity at the functional level (RKHS ball, etc …), or the $\ell_{1}$ norm, inducing sparsity at the level of the parameters $w$. Equivalently, (11) can be written as
@@
where $B$ depends on the regularization coefficient $\lambda$. The main two sources of error that need to be controlled in empirical risk minimization are the optimization and generalization error. The latter is handled using uniform convergence tools to control the convergence rate of the empirical risk towards the population one, when the parameters are constrained to the sublevel sets defined by the penalty. Note that we will not consider problems related to approximation error in this lecture.
@@
ensuring that the predictor $\hat{w}$ is a good proxy for the true minimizer. The standard way to achieve uniform control of the deviation between $\hat{L}(w)$, based on an available dataset $\mathcal{S}$, and $L(w)$ is to quantify the complexity of the hypothesis class $\mathcal{H}$ and regularity of the loss function, and to relate these quantities to the required sample complexity. A useful complexity measure for hypothesis classes is the Rademacher complexity, defined in its empirical form as
@@
$\sup_{h\in\mathcal{H}}|L(h)-\hat{L}(h)|\leq 2\mathcal{R}(l\circ\mathcal{H}\circ S)+c\sqrt{\frac{2/\delta}{m}}$ (16)
@@
$\mathbb{E}[\nabla l(\langle\bar{w}^{k},x_{i}\rangle,y_{i})]=\nabla L(\bar{w}^{k}).$
@@
Thus the error magnitude of the SGD and approximation error is the same. This means that we need to do at most $m$ iteration of SGD because when $m<T$, the dominant term in the previous bound becomes $O(\sqrt{\frac{B^{2}}{m}})$.
@@ 3.2 Stochastic optimization
A stochastic optimization problem is written as
$\min_{w\in{\mathcal{W}}}F(w)=\min_{w\in{\mathcal{W}}}\mathbb{E}_{z\sim{\mathcal{D}}}[f(w,z)]$ (20)
@@
$\min_{h}F(h)=\min_{h}\mathbb{E}_{z\sim{\mathcal{D}}}[f(h,z)],$
using the data samples $z_{i}\sim{\mathcal{D}}$, for $i=1,\ldots,m$, where $h$ is a function adapted to the particular model. Here is a short list of examples.
@@
    In both cases $L(w^{*})$ is the value of the cost function at the optimal point in the class: $L(w^{*})=\min_{\|w\|_{2}\leq B}L(w)$.
@@
Although we mentioned the effect of the step-size on the regularization in the direct approach, it is still not clear why we observe this phenomenon. Let us look back at the standard GD. The gradient descent minimizes the norm $\|w\|_{2}$. Indeed, at each step of the gradient descent we minimize its linear approximation given by the gradient $g^{(t)}=\nabla F(w^{(t)})$:
@@
The latter is linear function and thus its minimum is at infinity. Also, on the other hand, the linear approximation is not valid when we get far from the iterate $w^{(t)}$. To mitigate this effect, we can add a square regularizer $\|w-w^{t}\|_{2}^{2}/2\eta$, leading to
@@
In order to better understand this in the context of SGD, let us first introduce the notion of stability.
@@ 3.3 Stability
Define $\widehat{F}(w):=\frac{1}{m}\sum_{i=1}^{m}f_{i}(w)$. If $\tilde{w}$ is symmetric and $\beta(m)$ stable then
@@ 3.4 Strong convexity
The function $\Psi:{\mathcal{W}}\rightarrow\mathbb{R}$ is $\alpha$-strongly convex w.r.t. to a norm $\|w\|$ if for any $w,w^{\prime}\in\mathcal{W}$:
@@
Strong convexity essentially means that the function is bounded below by a quadratic function. This property depends on the norm unlike the convexity which can be defined on a vector space. Let us apply Eq.(23) for the optimum point $w_{0}:=\operatorname*{arg\,min}_{w}\Psi(w)$. Since $\nabla\Psi(w_{0})=0$ we get the following for every $w$:
@@
The latter means that the difference between the function value at any point and the optimal one is proportional to the distance from the optimal point. Let us now establish the connection between strong convexity and stability. For the dataset $S=\{z_{1},\ldots,z_{m}\}$ we define the regularized ERM (RERM) as follows:
@@
Here $\Psi$ is an $\alpha$-strongly convex function and $\widehat{F}$ is the empirical mean corresponding to the dataset $S$. We now recall the definition of Lipschitz continuity.
@@
One may notice that Lipschitz continuity implies some kind of “stability”, as it essentially means that small perturbation of the argument will not change the function value drastically. As for strong convexity, the Lipschitzness also depends on the norm of whatever normed space we are working on. This yields that $\|\nabla_{w}f(w,z)\|_{*}\leq G$, where $\|\cdot\|_{*}$ is the dual norm and $\nabla_{w}f(w,z)$ is viewed as an element of the dual of $\mathcal{W}$. We will assume that the norm is the same for both properties.
@@
In the rest of the lecture, without loss of generality, we may assume that the regularization function $\Psi$ is $1$-strongly convex. This is easily achieved by tuning the parameter $\lambda$ accordingly. Theorem 3.1 then yields the following
@@
where the last line is obtained by choosing $\lambda=\sqrt{\frac{2G^{2}}{\alpha m\sup_{w}\Psi(w)}}$. In particular, using the constraint $\Psi(w)\leq B$ and optimizing over the parameter $\lambda$, the last inequality can be rewritten as
@@
Let us now look back at the Risk minimization problem in the convex case:
@@
${\mathcal{W}}$ is assumed to be convex. If ${\sf loss}(\hat{y},y)$ is convex in $\hat{y}$, then the problem is convex. For a non-trivial loss, the composition $\textsf{loss}(h_{w}(x),y)$ is convex in $w$ only when $h_{w}(x)=\langle w,\phi(x)\rangle$. In this setting Lipschitz continuity may be established as follows. Assume that the loss function $\sf loss(y,y^{\prime})$ is $g$-Lipschitz continuous w.r.t. $y$. Then
@@
The online learning paradigm resembles the previously discussed stochastic optimization framework. The optimizer provides $w_{i}$. We give it to the adversary to compute $f(w_{i},z_{i})$ and then use the value of $f(w_{i},z_{i})$ to compute $w_{i+1}$.
@@
$\hat{w}_{m}(z_{1},\ldots,z_{m-1})=\arg\min_{w\in{\mathcal{W}}}\sum_{i=1}^{m-1}f(w,z_{i})+\lambda_{t}\Psi(w).$
This algorithm however, does not resemble the one-pass SGD algorithm (see also equation (22)):
@@
$\hat{w}_{m}^{\lambda}(z_{1},\ldots,z_{m-1})=\arg\min_{w\in{\mathcal{W}}}\frac{1}{m}\left\langle\sum_{i=1}^{m-1}\nabla f(w_{i},z_{i}),w\right\rangle+\lambda_{t}\Psi(w).$ (38)
@@ 3.5 Mirror descent
Figure 8: Bregman divergence.
@@
where $\Pi_{\Psi}^{{\mathcal{W}}}(w)=\min_{w^{\prime}\in{\mathcal{W}}}D_{\Psi}(w^{\prime}\mid w)$ is a projection step on $\mathcal{W}$ with respect to the Bregman distance. One may easily verify that if $\Psi$ is a quadratic function, then its Bregman divergence is also quadratic, and we recover the standard gradient descent algorithm, assuming the constraint set $\mathcal{W}$ is the domain of definition of $\Psi$. Indeed, taking $\Psi(x)=\frac{1}{2}\lVert x\rVert_{2}^{2}$, we obtain
@@
Figure 9: Suppose that the hypothesis class ${\mathcal{W}}$ is a subset of ${\mathcal{B}}$, where ${\mathcal{B}}$ is a Banach space. As we have seen above, the gradients live in the dual space ${\mathcal{B}}^{*}$. Then each iteration of the mirror descent consists of four steps. First we take the map $\nabla\Psi:{\mathcal{B}}\rightarrow{\mathcal{B}}^{*}$. Then we do a step in ${\mathcal{B}}^{*}$ in the direction of the gradient $\nabla f$. At the third step, we take $\Psi^{-1}$, which maps ${\mathcal{B}}^{*}$ to ${\mathcal{B}}$. The final step is the projection of this point to the hypothesis set ${\mathcal{W}}$.
@@
3.6 Mirror descent and implicit bias of optimization
We see that the generic formulation of the mirror descent algorithm allows to capture the geometry of an optimization problem through the Bregman distance of a chosen potential and the associated projection operator. For a given function $F(w)$ to optimize (over weights $w$), mirror descent implicitly minimizes an effective cost at each time step taking the form $\langle\nabla f(w_{t},z_{t}),w\rangle+\lambda_{t}D_{\Psi}(w\mid w_{t})$. We will see that, with some additional work, this framework can be used to understand implicit bias in optimization : for a given descent algorithm, we can find the potential that is being implicitly minimized and the optimality conditions that the fixed point of the descent method corresponds to. This will be the core topic of the next lecture.
@@ 4 Lecture 4 : Mirror descent and implicit bias of descent algorithms
We now have a good understanding of stochastic and online optimization for supervised learning. Consider an objective of the form:
@@
so that we have convexity with respect to $w\,.$ In what follows, we will abbreviate the pair $(x,y)$ as $z$. This is suitably generic because any data set is realizable in this form: We can choose $\phi$ as mapping $x$ to an indicator about its identity, and $w$ to select the appropriate label.
There are two reasons why studying convex optimization can be useful to understand deep learning: One is because it relates to mirror descent, and the second is that it allows us to discuss the geometry of the optimization problem and therefore the inductive bias.